# RAG-R1 : Incentivize the Search and Reasoning Capabilities of LLMs through Multi-query Parallelism

### v4

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across various tasks, while LLMs remain prone to generating hallucinated or outdated responses due to their static internal knowledge. Recent advancements in Retrieval-Augmented Generation (RAG) methods have aimed to enhance models’ search and reasoning capabilities through reinforcement learning (RL). Although these methods demonstrate promising results, they face challenges in training stability and encounter issues such as substantial inference time and restricted capabilities due to reliance on single-query mode. In this paper, we propose RAG-R1, a novel training framework designed to enable LLMs to adaptively leverage internal and external knowledge during the reasoning process. We further expand the generation and retrieval processes within the framework from single-query mode to multi-query parallelism, with the aim of reducing inference time and enhancing the model’s capabilities. Extensive experiments on seven question-answering benchmarks demonstrate that our method outperforms the strongest baseline by up to 13.2% and decreases inference time by 11.1%. The code and model checkpoints are available at https://github.com/inclusionAI/AgenticLearning/tree/main/RAG-R1.
@@ 1 Introduction
Large Language Models (LLMs) (Zhao et al., 2025; et al., 2025b) have demonstrated remarkable capabilities in various domains, including mathematical reasoning, question answering, and code generation. However, the knowledge encoded in these models is static, limiting their adaptability. As a result, LLMs are susceptible to producing hallucinated or outdated responses (Ji et al., 2023; Shuster et al., 2021; Asai et al., 2023) when dealing with complex or real-time issues, compromising their reliability. Therefore, it is essential to equip LLMs with access to external knowledge to ensure more accurate and grounded responses.
@@
Retrieval-Augmented Generation (RAG) (Gao et al., 2024; Fan et al., 2024) is a widely adopted approach to address this issue, broadening the model’s capability boundaries by integrating external knowledge into the generation process. Early efforts (Shao et al., 2023; Ram et al., 2023a) in this area focused on prompt-based strategies that guide LLMs through question decomposition, query rewriting and multi-turn retrieval. While effective, these approaches are limited by prompt engineering. Recent advances (Trivedi et al., 2023; Li et al., 2025; Asai et al., 2023; Guan et al., 2025; Wang et al., 2025) have increasingly emphasized the integration of reasoning capabilities within RAG. These approaches combine RAG with Chain of Thought (CoT) for step-by-step retrieval, or automatically generate intermediate retrieval chains that incorporate reasoning and employ Supervised Fine-Tuning (SFT) to enable learning of retrieval and reasoning. However, new insights (Chu et al., 2025) indicate that these methods may cause models to memorize solution paths, thus constraining their generalization to novel scenarios.
@@
Recently, reinforcement learning (RL) (Schulman et al., 2017) has demonstrated great potential in improving LLM performance by enhancing the reasoning capability. Reasoning models such as OpenAI-o1 (et al., 2024b) and Deepseek-R1 (et al., 2025b) indicate that utilizing RL with outcome-based rewards can enhance the model’s performance in mathematical and logical reasoning tasks. Within this paradigm, several works have explored enhancing the model’s search ability during reasoning through RL. R1-Searcher (Song et al., 2025) proposes a novel two-stage outcome-based RL approach designed to enhance the search capability of LLMs. Search-R1 (Jin et al., 2025) introduces a novel RL framework that enables LLMs to interact with search engines in an interleaved manner with their own reasoning. Despite significant improvements, these RL-based methods struggle with stable training due to the restricted capabilities of cold-start models (et al., 2025b). Furthermore, existing methods generate only a single search query whenever external retrieval is required, which presents two significant challenges: (1) Substantial Retrieval Iterations and Inference Time: The model generally requires multi-turn interleaved reasoning and search in single-query mode, particularly when dealing with multi-hop reasoning problems. This results in increased inference time, hindering its applicability in real-world scenarios. (2) Restricted External Knowledge: In single-query mode, the limited knowledge acquired from retrieval restricts the model’s ability to thoroughly explore the extent of its reasoning capability during training, thereby impacting its overall performance. We conducted a straightforward experiment based on Qwen2.5-72B-Instruct (et al., 2024a) following Jin et al. (2025) to validate the aforementioned challenges. We evaluated the model’s performance and average retrieval count when generating single query versus multiple queries in situations requiring retrieval. As illustrated in Figure 1, the multi-query method enhances both performance and inference efficiency compared to the single-query method, highlighting the limitations of the single-query mode.
@@
To address the aforementioned challenges, we propose RAG-R1, a novel training framework that enables LLMs to adaptively leverage internal and external knowledge during the reasoning process and improves the reasoning ability to answer questions correctly. We further expand the generation and retrieval processes within the framework from single-query mode to multi-query parallelism to reduce the model’s inference time and enhance its capabilities. Specifically, the training framework contains two stages, i.e., Format Learning Supervised Fine-Tuning and Retrieval-Augmented Reinforcement Learning. In the first stage, we thoughtfully generate samples that integrate reasoning and search to equip LLMs with the ability to adaptively leverage internal and external knowledge during the reasoning process and response in a think-then-search format. In the second stage, we employ outcome-based RL with a retrieval environment to improve the model’s ability to reason and dynamically retrieve external knowledge to answer questions correctly. Building upon the training framework, we expand the generation and retrieval processes from single-query mode to multi-query parallelism. This approach reduces retrieval rounds and inference time, supplying the model with more comprehensive and diverse information and ultimately boosting its performance.
@@
We conduct extensive experiments based on Qwen2.5-7B-Instruct (et al., 2024a) to verify the effectiveness of our proposed method. The results demonstrate its effectiveness, which achieves state-of-the-art performance on seven question answering benchmarks. In particular, our method utilizing multiple-query parallelism outperforms the strongest baseline by up to 13.2% and decreases inference time by 11.1%.
The contributions of this work are summarized as follows:
@@
We propose RAG-R1, a novel training framework that empowers LLMs with the ability to adaptively leverage internal and external knowledge during the reasoning process and improves the reasoning ability to answer questions correctly.
We further expand the generation and retrieval processes within the framework from single-query mode to multi-query parallelism, which not only reduces retrieval rounds and inference time of the model, but also further boosting its performance.
@@
Extensive experiments demonstrate the effectiveness of our proposed method, which achieves state-of-the-art performance on seven question answering benchmarks and significantly decreases retrieval rounds and inference time.
@@ 2 Related Work
Retrieval-Augmented Generation RAG enhances LLMs outputs by integrating external information with internal knowledge. Initial RAG methods employed static pipelines, coupling retrieved documents with LLMs via prompt engineering (Shi et al., 2023; Ram et al., 2023b; Lin et al., 2023; Korikov et al., 2024) to control generation. While effective for information-sourcing tasks, these approaches struggle with complex, multi-faceted queries requiring iterative retrieval. Subsequent research developed dynamic RAG frameworks (Jeong et al., 2024) that utilize LLM feedback for adaptive retrieval and generation control. Complementary efforts have focused on improving knowledge representation of retrieved documents (Edge et al., 2024; Guo et al., 2025) from augmentation perspectives. Most recently, advances (Jin et al., 2025; Li et al., 2025; Song et al., 2025; Sun et al., 2025; Wu et al., 2025) demonstrate significant scalability potential by training interaction trajectories to optimize LLM’s use of search engines for complex problem-solving queries. Following these advances, our method further enhances the interaction between LLMs and retrievers by leveraging LLM’s reasoning abilities.
Learning to Search Recent advances in Learning to Search have evolved from static query and generation template via SFT to dynamic and reward-driven search using RL. SFT-based approaches have demonstrated success in improving LLM capabilities in instruction following (Dong et al., 2024b; a; Li et al., 2024), robustness (Dong et al., 2024c), and domain-specific adaptation (Zhang et al., 2024). However, these methods are limited by a tendency to memorize solution paths, which constrains generalization and scalability. RL-based methods (et al., 2024b; 2025b; Shao et al., 2024; et al., 2025a) offer a promising alternative by enabling autonomous reasoning and decision-making. Recent works (Li et al., 2025; Song et al., 2025; Jin et al., 2025) further integrate LLMs into real-time search workflows, allowing interaction with live search engines. Although these methods support autonomous search engine invocation, the efficacy of search-as-a-tool remains underexplored. To address this, our approach introduces a two-stage training framework that enables parallel query generation, with the aim of retrieving more comprehensive and diverse document contexts.
@@ 3 Training Framework
In this section, we introduce the RAG-R1 training framework, which aims to empower LLMs with the ability to adaptively leverage internal and external knowledge during reasoning through two stages, i.e., Format Learning Supervised Fine-Tuning and Retrieval-Augmented Reinforcement Learning. Specifically, in the first stage, we thoughtfully generate samples that integrate reasoning and search, segmenting them into four categories. We then apply SFT based on these segmented samples to teach LLMs generating responses in a think-then-search format and enable it to leverage internal and external knowledge adaptively. In the second stage, we conduct data selection and employ outcome-based RL with a retrieval environment to enhance the model’s ability to reason and dynamically retrieve external knowledge to answer questions correctly. The overall framework is shown in Figure 2.
@@ 3.1 Format Learning Supervised Fine-Tuning
3.1.1 SFT Data Generation
To equip LLMs with the ability to adaptively leverage internal and external knowledge during reasoning and response in a think-then-search format, we first generate samples that integrate reasoning and search. Specifically, the system instruction guides the LLM to perform reasoning between think and /think whenever it receives new information. After completing the reasoning process, the LLM can generate search query between the designated search call tokens, search and /search, whenever external retrieval is necessary. The system will subsequently extract the search query and request an external search engine to retrieve relevant documents. The retrieved information is then appended to the existing sequence, enclosed within special retrieval tokens, information and /information, providing additional context for the next generation step. This process continues iteratively until reaching the maximum number of retrieval or the model generates a final answer enclosed within the special answer tokens, answer and /answer. The system instruction follows Jin et al. (2025) and is shown in Table 1.
@@ 3.1.1 SFT Data Generation
We use Qwen2.5-72B-Instruct as the generation model and employ a portion of the HotpotQA (Yang et al., 2018) training dataset for generation tasks.
3.1.2 SFT Data Segmentation and Training
After generating the samples, we selected those with correct answers and segment them into four categories according to specific segmentation points. As illustrated in Figure 2, we segment the samples at corresponding points whenever the model needs to perform reasoning or retrieval, thereby preventing the model from generating retrieval documents. In this manner, a complete sample might be split into multiple smaller samples, which can then be classified into four distinct categories.
By employing the first and second categories of samples, we train the model to respond by adaptively leveraging internal and external knowledge, respectively. In contrast, by utilizing the third and fourth categories of samples, we expect the model to perform reasoning and generate subsequent steps primarily based on external knowledge. Notably, the output section of the samples does not include retrieved documents, which aids in preventing the model from generating hallucinations.
After segmentation, we use these samples to perform SFT to teach the model generating responses in a think-then-search format and enable it to leverage internal and external knowledge adaptively. This stage aims to develop a highly capable model that can respond in a think-then-search format, serving as a cold-start model to enhance the stability of the subsequent RL training stage.
@@ 3.2 Retrieval-Augmented Reinforcement Learning
3.2.1 RL Data selection
After Format Learning SFT, we obtain a model that can adaptively leverage internal and external knowledge during the reasoning process and response in a think-then-search format. To further enhance the model’s reasoning ability and enable it to answer questions accurately, we begin with data selection to identify challenging yet answerable questions suitable for RL.
Specifically, we select the samples that generate incorrect answers in Section 3.1.1. The questions of these samples present a relative challenge to the model. We subsequently filter out samples that are inherently unanswerable due to incomplete data retrieval or model limitations. We implement stochastic sampling for Qwen2.5-72B-Instruct with the sampling temperature of 1.2 and the maximum number of retrieval to 10. For each question, we conduct up to 10 rollouts and retain only samples containing at least one correct solution. Finally, we obtain 2488 samples which are challenging yet answerable and randomly selected 25% of the correctly answered samples in Section 3.1.1 to construct the final training dataset.
3.2.2 RL with Retrieval
@@
where and represent the policy model and the reference model, respectively, both of which are initialized from the SFT model, is the reward function and is KL-divergence. denote input samples drawn from the dataset , and represent the generated outputs which is sampled from the policy model and retrieved from the retrieval system . The rollout process follows the same procedure detailed in Section 3.1.1. We employ the Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm to optimize the policy model, which is widely used in reinforcement learning due to its efficiency and reliability.
Retrieval Masked Loss In our framework, the rollout sequence consists of both LLM-generated tokens and retrieved tokens from the external retrieval system. To prevent retrieved tokens from interfering with the inherent reasoning and generation capabilities of the model, we implement a masked loss for these tokens. By computing the policy gradient objective exclusively on the LLM-generated tokens and excluding the retrieved content from optimization, this approach stabilizes training while preserving the adaptability and benefits of retrieval-augmented generation.
Reward Modeling The reward function serves as the primary training signal, which decides the optimization direction of RL. Inspired by et al. (2025b), we adopt a rule-based reward system that includes final answer rewards to assess the accuracy of responses. For instance, in factual reasoning tasks, we adopt rule-based criteria such as exact string matching (EM) to evaluate correctness:
where represents the extracted final answer from response and denotes the ground truth answer. We avoid incorporating format rewards because the SFT model has already learned to respond in a think-then-search format. Furthermore, following et al. (2025b), we do not apply neural reward models to avoid reward hacking and additional computational cost.

### v5

# RAG-R1: Incentivizing the Search and Reasoning Capabilities of LLMs through Multi-query Parallelism

@@ Abstract
Large Language Models (LLMs), despite their remarkable capabilities, are prone to generating hallucinated or outdated content due to their static internal knowledge. While Retrieval-Augmented Generation (RAG) integrated with Reinforcement Learning (RL) offers a solution, these methods are fundamentally constrained by a single-query mode, leading to prohibitive latency and inherent brittleness. To overcome these limitations, we introduce RAG-R1, a novel two-stage training framework centered around multi-query parallelism. Our framework enables LLMs to adaptively leverage internal and external knowledge during the reasoning process while transitioning from the single-query mode to multi-query parallelism. This architectural shift bolsters reasoning robustness while significantly reducing inference latency. Extensive experiments on seven question-answering benchmarks confirm the superiority of our method, which outperforms the strongest baseline by up to 13.7% and decreases inference time by 11.1%.
Code — https://github.com/inclusionAI/AWorld-RL/tree/main/RAG-R1
@@ 1 Introduction
Large Language Models (LLMs) (Survey_of_LLMs; deepseek-r1) have demonstrated remarkable capabilities in various domains, including mathematical reasoning, question answering, and code generation. However, the knowledge encoded in these models is static, limiting their adaptability. As a result, LLMs are susceptible to producing hallucinated or outdated responses (Survey_of_Hallucination; RAG_Reduce_Hallucination; self-rag) when dealing with complex or real-time issues, compromising their reliability. Therefore, it is essential to equip LLMs with access to external knowledge to ensure more accurate and grounded responses.
@@
Retrieval-Augmented Generation (RAG) (RAG_A_Survey; RAG_Meeting_LLMs) is a widely adopted approach to address this issue, broadening the model’s capability boundaries by integrating external knowledge into the generation process. Early efforts (ITER-RETGEN; In_Context_RAG) in this area concentrated on prompt-based strategies that guide LLMs through question decomposition, query rewriting and multi-turn retrieval. While effective, these approaches are constrained by the limitations inherent in prompt engineering. Recent advances (IRCoT; Search-o1; self-rag; DeepRAG; CoRAG) have increasingly emphasized the integration of reasoning capabilities within RAG. These approaches combine RAG with Chain of Thought (CoT) for step-by-step retrieval, or automatically generate intermediate retrieval chains that incorporate reasoning and employ Supervised Fine-Tuning (SFT) to enable learning of retrieval and reasoning. However, new insights (SFT_Memorizes) indicate that these methods may cause models to memorize solution paths, thus constraining their generalization to novel scenarios.
@@
Recently, reinforcement learning (RL) (PPO) has demonstrated great potential in improving LLM performance by enhancing the reasoning capability. Reasoning models such as OpenAI-o1 (openai-o1) and Deepseek-R1 (deepseek-r1) indicate that utilizing RL with outcome-based rewards can enhance the model’s performance in mathematical and logical reasoning tasks. Within this paradigm, several works have explored enhancing the model’s search ability during reasoning through RL. R1-Searcher (R1-Searcher) proposes a novel two-stage outcome-based RL approach designed to enhance the search capability of LLMs. Search-R1 (Search-R1) introduces a novel RL framework that enables LLMs to interact with search engines in an interleaved manner with their own reasoning. Despite significant improvements, these RL-based methods struggle with stable training due to the restricted capabilities of cold-start models (deepseek-r1). Furthermore, existing methods generate only a single search query whenever external retrieval is required, which presents two significant challenges: (1) Prohibitive Latency from Serial Execution. The single-query architecture mandates a serial execution model for multi-hop reasoning, where each step must await the completion of the preceding retrieval-inference cycle. Consequently, latency accumulates with each turn, rendering the model impractical for latency-sensitive, real-world applications. (2) Inherent Brittleness and Knowledge Confinement. The single-query approach is inherently brittle, as its entire reasoning trajectory depends on the success of each sequential step. An early, suboptimal query can lock the model into an unrecoverable path—a critical failure mode. This single-threaded process confines the model to a narrow slice of external knowledge, leaving it unable to navigate user ambiguity, explore alternative reasoning paths, or recover from an initial misstep. This fragility ultimately compromises its reasoning robustness and final performance. We conducted a straightforward experiment based on Qwen2.5-72B-Instruct (qwen2.5) following Search-R1 to validate the aforementioned challenges. As illustrated in Figure 1, we evaluated the model’s performance and average retrieval count on two Multi-Hop Question Answering benchmarks: HotpotQA (yang2018hotpotqa) and 2WikiMultiHopQA (xanh2020_2wikimultihop), comparing the generation of a single query to multiple queries in scenarios necessitating retrieval. The findings indicate that the multi-query method enhances the model’s performance and decreases retrieval iterations compared to the single-query approach, highlighting the limitations inherent in the single-query mode.
@@
To address the aforementioned challenges, we propose RAG-R1, a novel training framework that enables LLMs to adaptively leverage internal and external knowledge during the reasoning process and enhances their reasoning capabilities. We further expand the single-query mode to multi-query parallelism, an architectural shift that directly addresses prohibitive latency and inherent brittleness, thereby bolstering reasoning robustness and reducing inference time. Specifically, the training framework contains two stages, i.e., Format Learning Supervised Fine-Tuning and Retrieval-Augmented Reinforcement Learning. In the first stage, we thoughtfully generate samples that integrate reasoning and search to equip LLMs with the ability to adaptively leverage internal and external knowledge during the reasoning process and respond in a think-then-search format. In the second stage, we employ outcome-based RL with a retrieval environment to enhance the model’s ability to reason effectively and dynamically access external knowledge for accurate question answering. Building upon the training framework, we transition from the single-query mode to multi-query parallelism. By minimizing serial retrieval iterations, this approach drastically reduces overall inference time. Simultaneously, it arms the model with comprehensive evidence from multiple perspectives, directly bolstering its robustness against reasoning failures.
(a) Exact Match score
(b) Retrieval Iterations
@@
We conduct extensive experiments based on Qwen2.5-7B-Instruct (qwen2.5) to verify the effectiveness of our proposed method. The results demonstrate its effectiveness, which achieves state-of-the-art performance on seven question-answering benchmarks. In particular, our method utilizing multiple-query parallelism outperforms the strongest baseline by up to 13.7% and decreases inference time by 11.1%. The contributions of this work are summarized as follows:
@@
We propose RAG-R1, a novel training framework that empowers LLMs to adaptively leverage internal and external knowledge during the reasoning process and enhances their reasoning capabilities.
We transition from the single-query mode to multi-query parallelism to directly address prohibitive latency and inherent brittleness. This architectural shift bolsters reasoning robustness while significantly reducing inference time.
@@
Extensive experiments demonstrate the effectiveness of our proposed method, which achieves state-of-the-art performance on seven question-answering benchmarks and significantly decreases retrieval counts and inference time.
@@ 2 Related Work
Retrieval-Augmented Generation RAG enhances LLMs outputs by integrating external information with internal knowledge. Initial RAG methods employed static pipelines, coupling retrieved documents with LLMs via prompt engineering (replug; ralm; query-decomposing; multi-aspect-query) to control generation. While effective for information-sourcing tasks, these approaches struggle with complex, multi-faceted queries requiring iterative retrieval. Subsequent research developed dynamic RAG frameworks (adaptive-rag) that utilize LLM feedback for adaptive retrieval and generation control. Complementary efforts have focused on improving knowledge representation of retrieved documents (graph-rag; light-rag) from augmentation perspectives. Most recently, advances (Search-R1; Search-o1; R1-Searcher; zerosearch; masksearch) demonstrate significant scalability potential by training interaction trajectories to optimize LLM’s use of search engines for complex problem-solving queries. Following these advances, our method further enhances the interaction between LLMs and retrievers by leveraging LLM’s reasoning abilities.
Learning to Search Recent advances in Learning to Search have evolved from static query and generation template via SFT to dynamic and reward-driven search using RL. SFT-based approaches have demonstrated success in improving LLM capabilities in instruction following (vif-rag; self-play; retrollm), robustness (dpa-rag), and domain-specific adaptation (raft). However, these methods are limited by a tendency to memorize solution paths, which constrains generalization and scalability. RL-based methods (openai-o1; deepseek-r1; grpo; qwen3) offer a promising alternative by enabling autonomous reasoning and decision-making. Recent works (Search-o1; R1-Searcher; Search-R1) further integrate LLMs into real-time search workflows, allowing interaction with live search engines. Although these methods support autonomous search engine invocation, the efficacy of search-as-a-tool remains underexplored. To address this, our approach introduces a two-stage training framework that enables parallel query generation, with the aim of retrieving more comprehensive and diverse document contexts.
@@ 3 Training Framework
In this section, we introduce the RAG-R1 training framework, which aims to empower LLMs with the ability to adaptively leverage internal and external knowledge during reasoning through two stages, i.e., Format Learning Supervised Fine-Tuning and Retrieval-Augmented Reinforcement Learning. Specifically, in the first stage, we thoughtfully generate samples that integrate reasoning and search, segmenting them into four categories. We then apply SFT based on these segmented samples to equip LLMs with the ability to generate responses in a think-then-search format and leverage internal and external knowledge adaptively. In the second stage, we conduct data selection and employ outcome-based RL with a retrieval environment to enhance the model’s ability to reason and dynamically retrieve external knowledge to answer questions correctly. The overall framework is shown in Figure 2.
@@ 3.1 Format Learning Supervised Fine-Tuning
SFT Data Generation
To equip LLMs with the ability to adaptively leverage internal and external knowledge during reasoning and respond in a think-then-search format, we first generate samples that integrate reasoning and search. Specifically, the system instruction guides the LLM to perform reasoning between think and /think whenever it receives new information. After completing the reasoning process, the LLM can generate search query between the designated search call tokens, search and /search, whenever external retrieval is necessary. The system will subsequently extract the search query and request an external search engine to retrieve relevant documents. The retrieved information is then appended to the existing sequence, enclosed within special retrieval tokens, information and /information, providing additional context for the next generation step. This process continues iteratively until reaching the maximum number of retrieval or the model generates a final answer enclosed within the special answer tokens, answer and /answer. The system instruction follows Search-R1 and is shown in Table 1.
@@
We use Qwen2.5-72B-Instruct as the generation model and employ a portion of the HotpotQA (yang2018hotpotqa) training dataset for generation tasks.
SFT Data Segmentation and Training
After generating the samples, we selected those with correct answers and segment them into four categories according to specific segmentation points. As illustrated in Figure 2, we segment the samples at corresponding points whenever the model needs to perform reasoning or retrieval, thereby preventing the model from generating retrieval documents. In this manner, a complete sample might be split into multiple smaller samples, which can then be classified into four distinct categories. More details can be found in Appendix B.
By employing the first and second categories of samples, we train the model to respond by adaptively leveraging internal and external knowledge, respectively. In contrast, by utilizing the third and fourth categories of samples, we expect the model to perform reasoning and generate subsequent steps primarily based on external knowledge. Notably, the output section of the samples is designed to exclude retrieved documents, which aids in preventing the model from generating hallucinations.
After segmentation, we apply SFT to these samples to train the model to generate responses in a think-then-search format and leverage internal and external knowledge adaptively. This stage aims to develop a highly capable model that can respond in a think-then-search format, serving as a cold-start model to enhance the stability of the subsequent RL training stage.
@@ 3.2 Retrieval-Augmented Reinforcement Learning
RL Data selection
After Format Learning SFT, we obtain a model that can adaptively leverage internal and external knowledge during the reasoning process and respond in a think-then-search format. To further enhance the model’s reasoning ability and enable it to answer questions accurately, we begin with data selection to identify challenging yet answerable questions suitable for RL.
Specifically, we select the samples that generate incorrect answers in Section 3.1. The questions of these samples present a relative challenge to the model. We subsequently filter out samples that are inherently unanswerable due to incomplete data retrieval or model limitations. We implement stochastic sampling for Qwen2.5-72B-Instruct with the sampling temperature of 1.2 and the maximum number of retrieval to 10. For each question, we conduct up to 10 rollouts and retain only samples containing at least one correct solution. Finally, we obtain 2488 samples which are challenging yet answerable and randomly selected 25% of the correctly answered samples in Section 3.1 to construct the final training dataset.
RL with Retrieval
@@
where and represent the policy model and the reference model, respectively, both of which are initialized from the SFT model, is the reward function and is KL-divergence. denote input samples drawn from the dataset , and represent the generated outputs which is sampled from the policy model and retrieved from the retrieval system . The rollout process follows the same procedure detailed in Section 3.1. We employ the Proximal Policy Optimization (PPO) (PPO) algorithm to optimize the policy model, which is widely used in reinforcement learning due to its efficiency and reliability.
Retrieval Masked Loss. In our framework, the rollout sequence consists of both LLM-generated tokens and retrieved tokens from the external retrieval system. To prevent retrieved tokens from interfering with the inherent reasoning and generation capabilities of the model, we implement a masked loss for these tokens. By computing the policy gradient objective exclusively on the LLM-generated tokens and excluding the retrieved content from optimization, this approach stabilizes training while preserving the adaptability and benefits of retrieval-augmented generation.
Reward Modeling. The reward function serves as the primary training signal, which decides the optimization direction of RL. Inspired by deepseek-r1, we adopt a rule-based reward system that includes final answer rewards to assess the accuracy of responses. For instance, in factual reasoning tasks, we adopt rule-based criteria such as exact string matching (EM) to evaluate correctness:
where represents the extracted final answer from response and denotes the ground truth answer. This method focuses on ensuring that the final answer matches the expected truth precisely. We avoid incorporating format rewards because the SFT model has already learned to respond in a think-then-search format. This strategy enables us to concentrate on the factual accuracy of the responses, leveraging the inherent structured approach of the model without imposing additional format-based conditions. Furthermore, following deepseek-r1, we do not apply neural reward models to avoid reward hacking and additional computational cost.